solve each system using the substitution method.
x=y-5
2x+5y=4

Answer by london maths tutor(243)

x=y-5 ... equation 1
2x+5y=4 ... equation 2
Put equation 1 into equation 2
2(y-5) + 5y = 4
2y - 10 + 5y = 4
7y - 10 = 4
7y = 4 + 10
7y = 14
y = 14/7
y = 2
Sub y = 2 into equation 1
x = 2 - 5
x = -3
Answer : x = -3 and y = 2